OpenStudy (anonymous):

is 2 1/4 + 1.5 - 1 1/2 < ,> 3 5/8 - 1.46?

OpenStudy (anonymous):

show work plz

OpenStudy (agreene):

2 1/4 + 1.5 - 1 1/2 = 9/4 or 4 1/4 or 2.25 if you prefer 3 5/8 - 1.46 = 433/200 or 2 33/200 or 2.165 2.25 > 2.165 therefore 2 1/4 + 1.5 - 1 1/2 > 3 5/8 - 1.46

OpenStudy (anonymous):

I couldn't change 3 5/8 to a decimal...when I worked it